### 14 Mar 2024 v3

# The set of $\Phi$ badly approximable matrices has full Hausdorff dimension

Abstract.
Recently Koivusalo, Levesley, Ward and Zhang introduced the set of simultaneously $\Phi$-badly approximable real vectors of $\mathbb{R}^{m}$ with respect to an approximation function $\Phi$, and determined its Hausdorff dimension for the special class of power functions $\Phi(t)=t^{-\tau}$. We refine this by naturally extending the formula to arbitrary decreasing functions, in terms of the lower order of $1/\Phi$ at infinity. We also provide an alternative, rather mild condition on $\Phi$ for this conclusion. Moreover, our results apply in the general matrix setting, and we establish an according formula for packing dimension as well. Thereby we also complement a recent refinement by Bandi and de Saxcé on the smaller set of exact approximation with respect to $\Phi$. Our basic tool is (a uniform variant of) the variational principle by Das, Fishman, Simmons, Urbański. We also prove some new lower estimates regarding the set of exact approximation order in the matrix setting, which are sharp in special instances. For this we combine the result by Bandi and de Saxcé with a method developed by Moshchevitin.
Middle East Technical University, Northern Cyprus Campus, Kalkanli, Güzelyurt
johannes@metu.edu.tr ; jschleischitz@outlook.com
Keywords: Hausdorff dimension, packing dimension, exact approximation, variational principle
Math Subject Classification 2020: 11H06, 11J13
1. Introduction: Sets of exact and bad approximation
Let $m,n$ be positive integers and $\Phi:\mathbb{N}\to(0,\infty)$ be a function with the property that
(1) $\Phi(t)\leq t^{-n/m},\qquad t\geq 1.$
Note that we do not assume monotonicity. Define the set of $\Phi$-approximable matrices $W(\Phi)$ to be the set of real $m\times n$ matrices $A$ such that
$\|A\textbf{q}-\textbf{p}\|\leq\Phi(\|\textbf{q}\|)$
[…]
$\tau=\liminf_{t\to\infty}-\frac{\log\Phi(t)}{\log t}\in[\frac{n}{m},\infty].$
The most natural instances are power functions $\Phi(t)=t^{-\tau}$ with $\tau\geq n/m$, where the right hand side is constant. If $\tau$ is finite, in this case for simplicity let $W_{\tau}=W(t\to t^{-\tau})$, which is essentially the set of matrices approximable to order $\tau$. For $\tau=\infty$, accordingly let $W_{\infty}=\cap_{\tau\in\mathbb{R}}W_{\tau}$ be the set of $m\times n$ Liouville matrices. Notice that if $\tau=n/m$, the set $W_{\tau}$ becomes the entire space of real $m\times n$ matrices by a well-known variant of Dirichlet’s Theorem. This also explains why our restriction (1) is natural, although strictly speaking the set $W(\Phi)$ does not necessarily consist of all real $m\times n$ matrices if $\tau<n/m$ for a general $\Phi$ (there are counterexamples if the upper order of $1/\Phi$ at infinity exceeds $n/m$ and $\Phi$ fluctuates fast; it is however true for any monotonic $\Phi$). Define the set
$Exact(\Phi)=W(\Phi)\setminus\bigcup_{c>1}W(\Phi/c)$
of “exactly” $\Phi$ approximable $m\times n$ matrices. Indeed, it consists of the matrices that are $\Phi$-approximable, but not $\Phi/c$-approximable no matter how close $c>1$ is to $1$. It is suggestive to believe in the following
Conjecture 1.
The set $Exact(\Phi)$ has the same Hausdorff dimension as $W_{\tau}$ as soon as
(2) $\Phi(t)=o(t^{-n/m}),\qquad t\to\infty.$
We will occasionally for brevity refer to the dimension as “full” in this case. For $m=n=1$ the conjecture is essentially the content of [5, Problems 1 & 2]. Possibly (2) can be relaxed in some ways, however (1) is insufficient at least for $m=n=1$ due to well-known facts about the Lagrange spectrum, as already pointed out in [5]. For $m=n=1$ and decreasing $\Phi$, in fact under a certain weaker assumption ()($\ast$)( ∗ ) on $\Phi$, the conjecture was verified by Bugeaud and Moreira [5, Theorem 3], see also the preceding work by Bugeaud [3, 4], both dimensions being $2/(\tau+1)$. Only very recently, during the process of finishing the present paper, this was extended to the case of simultaneous approximation to $m$ real numbers, i.e. $n=1$, by Bandi and de Saxcé [1], the full dimension being $(m+1)/(\tau+1)$. The latter paper [1] again imposed $\Phi$ to be decreasing and satisfying (2). While it is likely that the method in [1] can be extended to the general matrix setting by similar arguments, in all other cases of $m,n$ the conjecture remains yet unknown even assuming monotonic $\Phi$. The case of non-monotonic $\Phi$ is not completely understood even for $m=n=1$.
Prior to [1], in preprint [8] from autumn 2023 as well, it is shown that again for $n=1$ and the special case $\Phi(t)=t^{-\tau}$, $\tau>1/m$, the larger set of $\Phi$-badly approximable vectors defined by
$Bad(\Phi)=\bigcup_{c>1}W(\Phi)\setminus W(\Phi/c)=W(\Phi)\setminus\bigcap_{c>1}W(\Phi/c)$
has full Hausdorff dimension $(m+1)/(\tau+1)$, a weaker result. The latter paper [8] in turn improved on results from [2]. For any $\Phi$ we have the obvious inclusions
(3) $Exact(\Phi)\subseteq Bad(\Phi)\subseteq\bigcap_{\epsilon>0}W_{\tau-\epsilon}.$
The methods from [1, 8] are rather different. We refer to [1, 5, 8] for a more comprehensive history of exact approximation and more relevant literature. We want to add that in [10, Corollary 7] it is shown that again for $n=1$, for arbitrary $\Phi$ with $\tau>\tau_{m}$ large enough, the packing dimension of $Exact(\Phi)$ equals $m$, hence is full literally. Hereby suitable $\tau_{m}$ can be explicitly determined, decreasing to the limit $(3+\sqrt{5})/2=2.6180\ldots$ with initial term $\tau_{1}=(5+\sqrt{17})/2=4.5615\ldots$. It is natural to believe in an analogue of Conjecture 1 for packing dimension. The method from [10] is very different from any other previous work on the topic, including the present paper.
Acknowledgements: The paper was started during a visit for an International Visitor Program at the University of Sydney. The author thanks the University of Sydney for the generous support.
2. New results
In this paper, we show that under some rather mild additional regularity assumption on $\Phi$, the Hausdorff dimension of $Bad(\Phi)$ is full, i.e. the same as for $W_{\tau}$. Thereby we properly extend the result in [8] and complement [1]. Moreover, we obtain an analogous result for packing dimension. As an alternative to assuming $\Phi$ monotonic, a widely used simplifying condition used in [1, 8], the following rather mild condition (C1) suffices: For any $c>1$, there exists $d=d(c)>1$ such that
(C1) $c^{-1}t\leq\tilde{t}\leq ct\quad\Longrightarrow\quad d^{-1}\Phi(t)\leq\Phi(\tilde{t})\leq d\Phi(t),$
for any positive integers $t,\tilde{t}$. We now turn towards our main result, a discussion on condition (C1) is given below. Denote by $\dim_{H}$ and $\dim_{P}$ the Hausdorff and packing dimensions of a set respectively. We show
Theorem 2.1.
Let $\Phi$ satisfy (1) and additionally assume that either condition (C1) holds or $\Phi$ is decreasing. Then
(4) $\dim_{H}(Bad(\Phi))=\dim_{H}(W_{\tau})=(n-1)m+\frac{m+n}{1+\tau},$
and
(5) $\dim_{P}(Bad(\Phi))=\dim_{P}(W_{\tau})=mn.$
An alternative sufficient condition for (5), applicable for the smaller sets $Exact(\Phi)$, will be given in Theorem 2.3 below. The dimensions of $W_{\tau}$ in the theorem are well-known, see [8] for a reference. For the more interesting identities involving $Bad(\Phi)$, the upper estimates are rather obvious, in view of the right inclusion of (3) it suffices to take the limit of the (continuous) dimension formulas for the sets $W_{\tau-\epsilon}$ as $\epsilon\to 0$. As in [1, 8], the reverse lower bounds are the substantial result.
We discuss (C1) in more detail now. It can be readily checked that we only need to assume the implication holds for some $c>1$. Moreover, if (C1) holds for some pair $c_{0},d_{0}=c_{0}^{\rho_{0}}$, via partitioning $(1,\infty)$ into intervals $(1,c_{0}],[c_{0},c_{0}^{2}),[c_{0}^{2},c_{0}^{4}),[c_{0}^{4},c_{0}^{8}),\ldots$, it can be seen that for any $c>1$ it holds for $d=\max\{d_{0},c^{2\rho_{0}}\}$. In particular if (C1) holds then we can assume $\rho(c)=\log d/\log c\ll_{\Phi,\epsilon}1$ bounded on any interval $c\in[1+\epsilon,\infty)$. Essentially, condition (C1) says that $\Phi$ does not deviate too fast on short intervals. It is in particular true for all power functions $t\to t^{-\tau}$, as treated in [8]. It is easily seen that the assumptions $\Phi$ being monotonic and (C1) are independent. Moreover (C1) it is independent from the condition ()($\ast$)( ∗ ) from [5, § 2] for $m=n=1$ already mentioned in § 1, which can be stated as $\Phi(ct)\leq 4c^{\rho}\Phi(t)$ for any $c>1$ and some absolute $\rho\in\mathbb{R}$, uniformly in $t\geq t_{0}$. Hence $\Phi$ may decay arbitrarily fast, so for example $\Phi(t)=e^{-t}$ satisfies ()($\ast$)( ∗ ) but not (C1). On the other hand, for example the piecewise defined function $\Phi(t)=2^{N}t^{-3}$ for $t\in[2^{N},2^{N+1})$, $N\in\mathbb{N}$, satisfies (C1) but not ()($\ast$)( ∗ ). Note that $\rho(c)\to\infty$ as $c\to 1^{+}$ for the latter.
We compare Theorem 2.1 with the claims from [1, 5, 8, 10] discussed in § 1. Claim (4) refines results from [1, 8] in two aspects: Firstly, we can deal with the general matrix setup instead of just simultaneous approximation $n=1$. Secondly, our result applies to a wider class of approximation functions that are possibly not decreasing (especially [8] restricted to $t\to t^{-\tau}$, $\tau>1/m$). Our above comparison of (C1) versus condition ()($\ast$)( ∗ ) from [5] shows that even for $m=n=1$ there is some new information derived from (4). In particular for $n=1$ and decreasing $\Phi$ formula (4) is implied by the stronger claims from [1] (however we give an independent proof). It may however be viewed as a disadvantage that our proof relies on the heavy machinery in [6]. Also, as an artifact of these results in [6], we cannot provide a fixed constant $c>1$ such that the full dimension claim holds for the set $W(\Phi)\setminus W(\Phi/c)\subseteq Bad(\Phi)$, which is possible in [8]. The packing dimension result (5) can be regarded a variant of [10, Corollary 7]. Advantages are that it applies for matrix setting and only the weaker natural bound $\tau\geq n/m$ on $\tau(\Phi)$ is needed. However, we must assume (C1) or monotonic $\Phi$ instead and cannot provide any information on the smaller sets $Exact(\Phi)$ treated in [10]. In terms of proofs, our strategy below will use similar arguments as in [1], also based on parametric geometry of numbers (template theory), however with some twists. Most notably, we directly apply the variational principle from [6], in place of the Cantor set construction in [1] that allows for treating the smaller set $Exact(\Phi)$.
In fact our proof of Theorem 2.1 below will show the following refined claim. As usual $a\asymp_{.}b$ means $a\ll_{.}b\ll_{.}a$, where $a\ll_{.}b$ in turn means $a\leq cb$ for some constant $c>0$ depending on the subscript variables only, while $a,b$ may depend on other parameters as well. If there are no subscript variables the constants are absolute.
Theorem 2.2.
Let $\Phi$ be as in Theorem 2.1. Let $\chi:\mathbb{N}\to\mathbb{R}$ by any function/sequence (increasing arbitrarily fast). Then the metrical conclusions of Theorem 2.1 hold for the set of real $m\times n$ matrices $A$ for which the following two properties hold:
  • There exists a sequence $(t_{k}^{\prime})_{k\geq 1}$ of integers, depending on $A$, such that
    $t_{k}^{\prime}>\chi(t_{k}^{\prime}),$
    and
    $\Phi(t_{k}^{\prime})=t_{k}^{\prime-\tau+o(1)},\qquad k\to\infty,$
    and
    $\|\textbf{q}_{k}\|=t_{k}^{\prime},\qquad\Phi(t_{k}^{\prime})\ll_{A}\|A\textbf{q}_{k}-\textbf{p}_{k}\|<\Phi(t_{k}^{\prime}),\qquad k\geq 1,$
    holds for some integer vector $(\textbf{p}_{k},\textbf{q}_{k})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\boldsymbol{0}\})$.
  • For any $(\textbf{p},\textbf{q})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\boldsymbol{0}\})$ not among the $(\textbf{p}_{k},\textbf{q}_{k})$ we have
    $\|A\textbf{q}-\textbf{p}\|\gg_{A}\|\textbf{q}\|^{-n/m},\qquad k\geq 1.$
In view of (1) this clearly implies Theorem 2.1. In order to apply [6], the implied constants in the theorem will tend to infinity to derive the full dimension results. It may be true that analogous extensions of the results stated in [1] can be derived from their method as well, as the proofs show substantial similarities.
Finally, we want to state some new claims for the sets $Exact(\Phi)$. From [1, 10] when combined with a method originating in [9], under certain conditions on $\Phi$ we derive a non-trivial lower bound for the Hausdorff dimension and full packing dimension in the matrix setting. Let
$\theta=\theta_{m,n}:=\frac{\max\{n+1,m+2\}}{m}>1.$
Then our result reads as follows.
Theorem 2.3.
The following claims hold:
  • (a)
    Let $\Phi$ induce $\tau>\theta$ and be decreasing. Then
    (6) $\dim_{H}(Exact(\Phi))\geq m(n-1)+\frac{m+1}{1+\tau}.$
  • (b)
    Let $\Phi$ induce $\tau>\max\{\theta,\frac{5+\sqrt{17}}{2}\}$ (not necessarily decreasing). Then
    (7) $\dim_{H}(Exact(\Phi))\geq m(n-1)$
    and
    (8) $\dim_{P}(Exact(\Phi))=mn.$
Remark 1.
The weaker property ()($\ast$)( ∗ ) from [5] discussed in § 1 instead of $\Phi$ decreasing, together with $\tau>\theta$, suffice to settle the intermediate bound
$\dim_{H}(Exact(\Phi))\geq m(n-1)+\frac{2}{\tau+1},$
in between (6), (7). In particular for $\tau=\infty$ we get full dimension, refining (4) from Theorem 2.1 in this case. See § 6.2 below for a proof.
For $\tau=\infty$, the Hausdorff dimension in (6) and (7) is full, a refinement of Theorem 2.1. The condition $\tau>\theta>n/m$ is more restrictive than (1) and (2). It can be slightly relaxed as the proof of Lemma 5.2 below shows, however we cannot decrease the bound for $\tau$ with our method, in particular (2) is insufficient in both claims. For $n=1$ estimate (6) is implied by [1]. The value $(5+\sqrt{17})/2$ in the condition for (7), (8) derived from [10] can be relaxed if $m>1$.
3. Combined graphs, templates and the variational principle
For $1\leq j\leq m+n$ and $A\in\mathbb{R}^{m\times n}$, and a parameter $Q\geq 1$, let $\lambda_{j}(Q)$ be the $j$-th successive minimum with respect to the parametric convex body
$K(Q)=[-Q,Q]^{n}\times[-Q^{-n/m},Q^{-n/m}]^{m}\subseteq\mathbb{R}^{m+n}$
and the lattice
$\{(\textbf{q},A\textbf{q}-\textbf{p}):\;\textbf{q}=(q_{1},\ldots,q_{n})\in\mathbb{Z}^{n},\;\textbf{p}=(p_{1},\ldots,p_{m})\in\mathbb{Z}^{m}\}\subseteq\mathbb{R}^{m+n}.$
Further derive continuous piecewise linear functions
$h_{j}(q)=\log(\lambda_{j}(e^{q})),\qquad 1\leq j\leq m+n,\;q>0.$
Thereby every real $m\times n$ matrix $A$ gives rise to successive minima functions $\textbf{h}=(h_{1},\ldots,h_{m+n})$ on $(0,\infty)$. We call the union of the graphs $(q,h_{j}(q))$, $q>0$ over $1\leq j\leq m+n$ in $\mathbb{R}^{2}$ the combined graph associated to $A$. The above definition is formally slightly different than the formulation using homogeneous dynamics from [6]. We rather follow the setup of Schmidt and Summerer [12] but for general $m,n$, however the two formalisms are easily verified to be equivalent. Define the trajectory associated to $(\textbf{p},\textbf{q})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\textbf{0}\})$ as
(9) $T_{\textbf{p},\textbf{q}}(q)=\max\left\{\log\|\textbf{q}\|-\frac{1}{n}q\;,\;\log\|A\textbf{q}-\textbf{p}\|+\frac{1}{m}q\right\}.$
Then $h_{1}(q)$ is the minimum of all trajectories
$h_{1}(q)=\min_{\textbf{p},\textbf{q}}T_{\textbf{p},\textbf{q}}(q)=\min_{\textbf{p},\textbf{q}}\max\left\{\log\|\textbf{q}\|-\frac{1}{n}q\;,\;\log\|A\textbf{q}-\textbf{p}\|+\frac{1}{m}q\right\}$
with minimum taken over all integer vectors $(\textbf{p},\textbf{q})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\textbf{0}\})$. More generally, for $1\leq j\leq m+n$, we have
$h_{j}(q)=\min\max_{1\leq i\leq j}T_{\textbf{p}_{i},\textbf{q}_{i}}(q)$
with the minimum taken over all tuples of $j$ linearly independent integer vectors $(\textbf{p}_{i},\textbf{q}_{i})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\textbf{0}\})$. Since any $h_{j}$ locally coincides with some trajectory, it is clear that it is indeed piecewise linear with slopes among $\{-1/n,1/m\}$.
We next introduce templates. Following [6, Definition 4.1], an $m\times n$-template consists of piecewise linear functions $\textbf{f}(q)=(f_{1}(q),\ldots,f_{m+n}(q))$ with a certain finite set of slopes, including $\{-1/n,1/m\}$ as smallest and largest elements, and several other restrictions, we omit stating the details here for brevity. We highlight that the sum of the $f_{j}(q)$, thus also of the derivatives where defined, vanishes on $q\in(0,\infty)$.
Each template gives rise to a local contraction rate $\delta(q)\in[0,mn]$ for each $q>0$ where f is differentiable, depending on the slope vector $\textbf{f}^{\prime}(q)$ at $q$, again we omit details here to be found in [6]. The lower and upper limits
[…]
are called lower and upper contraction rate of the template. Denote $\mathcal{T}_{m,n}$ the set of $m\times n$ templates.
Definition 1.
We say a combined graph h is $T>0$ close to a template $\textbf{f}\in\mathcal{T}_{m,n}$ if
$\sup_{q>0}\max_{1\leq j\leq m+n}|f_{j}(q)-h_{j}(q)|\leq T.$
We identify $A\in\mathbb{R}^{m\times n}\cong\mathbb{R}^{mn}$ giving rise to the natural $mn$-dimensional Lebesgue measure on the sets of $m\times n$ matrices. A partial claim of the landmark result [6, Theorem 4.7] can be formulated as follows:
Theorem 3.1 (Das, Fishman, Simmons, Urbański).
Let $\textbf{f}\in\mathcal{T}_{m,n}$ is a template with lower and upper contraction rates $\underline{\delta}(\textbf{f})$ and $\overline{\delta}(\textbf{f})$. Given $\varepsilon>0$, there is some absolute $T=T(\varepsilon)>0$ independent of f so that the set of real matrices $A\in\mathbb{R}^{m\times n}$ inducing a combined graph $\textbf{h}(A)$ that is $T$ close to f has Hausdorff dimension at least $\underline{\delta}(\textbf{f})-\varepsilon$ and packing dimension at least $\overline{\delta}(\textbf{f})-\varepsilon$.
4. Proof of Theorem 2.1 upon assumption (C1)
4.1. Construction of a template
Let $\Phi$ satisfy (1), (C1). We first construct a template $\textbf{f}=\textbf{f}(\Phi)$, in a similar way to the proof of [6, Theorem 3.17]. Take a lacunary sequence of integers $t_{k}>1$ such that
(10) $\lim_{k\to\infty}-\frac{\log\Phi(t_{k})}{\log t_{k}}=\tau,$
where by lacunary we mean
(11) $\lim_{k\to\infty}\frac{\log t_{k+1}}{\log t_{k}}=\infty.$
Note $\tau=\infty$ is possible. We restrict to the class $\mathcal{C}\subseteq\mathcal{F}_{m,n}$ of templates $\textbf{f}=(f_{1},\ldots,f_{m+n})$ with the property
$f_{2}(q)=f_{3}(q)=\cdots=f_{m+n}(q)=-\frac{f_{1}(q)}{m+n-1},\qquad q>0.$
Moreover we exclude the template with all component functions constant 00 for all $q\geq q_{0}$. Then for each such template $\textbf{f}\in\mathcal{C}$, the first component $f_{1}$ has slopes among $\{-1/n,0,1/m\}$, and we have
$f_{1}(q)=f_{2}(q)=\cdots=f_{m+n}(q)=0$
for $q$ on a sequence of pairwise disjoint closed intervals $I_{1}=[a_{1},b_{1}],I_{2}=[a_{2},b_{2}],\ldots$ on the first axis, where $a_{1}<b_{1}<a_{2}<b_{2}<\ldots$.
For simplicity let $c_{k}=a_{k+1}$ so that $J_{k}=(b_{k},c_{k})$ denotes the open interval between $I_{k}$ and $I_{k+1}$. Then in each $J_{k}$, the function $f_{1}$ has a unique local minimum, say at $q_{k}\in J_{k}$, and $f_{1}$ decreases with slope $-1/n$ on $(b_{k},q_{k})$ and increases with slope $1/m$ on the intervals $(q_{k},c_{k})$. The remaining $f_{j}$, $2\leq j\leq m+n$, have accordingly slopes $n^{-1}/(m+n-1)$ and $-m^{-1}/(m+n-1)$ on the respective intervals. We choose our concrete template $\textbf{f}\in\mathcal{C}$ so that these local minima of $f_{1}$ are at positions $(q_{k},f_{1}(q_{k}))$ with
(12) $q_{k}=(\log t_{k}-\log\Phi(t_{k}))\frac{1}{m^{-1}+n^{-1}},\qquad f_{1}(q_{k})=\log t_{k}-\frac{1}{n}q_{k}-\log\Delta,$
for $\Delta\geq 1$ to be chosen later. This resembles the minimum of a fictive trajectory $T_{\textbf{p},\textbf{q}}$ with $\|\textbf{q}\|=t_{k}$ and $\|A\textbf{q}-\textbf{p}\|=\Phi(\|\textbf{q}\|)/\Delta$. It is easily verified that the $q_{k}$, and thus the $t_{k}$, uniquely determine the template and
(13) $b_{k}=q_{k}+nf_{1}(q_{k}),\qquad c_{k}=q_{k}-mf_{1}(q_{k})$
hold for $k\geq 1$. Recall that $f_{1}(q_{k})<0$, so indeed $b_{k}<c_{k}$.
[…]
Figure 1: Sketch of template with slopes
We claim two statements that when combined almost directly yield the theorem in the version where we assume (C1). Recall the contraction rates of a template from the last section.
Lemma 4.1.
Let $\Phi$ satisfy (1), (C1) and $T>0$. Any matrix $A$ whose induced combined graph $\textbf{h}=\textbf{h}_{A}$ is $T$ close to f constructed above belongs to $Bad(\Phi)$.
Lemma 4.2.
The template f above has lower and upper average contraction rate
$\underline{\delta}(\textbf{f})=(n-1)m+\frac{m+n}{1+\tau},\qquad\underline{\delta}(\textbf{f})=mn,$
thus via Theorem 3.1 induces a set of matrices with combined graph $O(1)$ close to it with the desired Hausdorff and packing dimensions.
The latter lemma is again proved very similarly to [6, Theorem 3.17] so there is not much originality to it, the proof of the former lemma is however more involved. We prove the lemmas below.
4.2. Proof of Lemma 4.1
The claim follows from the next geometrically very intuitive observation, whose exact proof however turns out rather lengthy.
Lemma 4.3.
Let $\textbf{f}\in\mathcal{T}_{m,n}$ be the template constructed in §4.1. Let $T>0$ be a parameter and h be the combined graph associated to some $A\in\mathbb{R}^{m\times n}$ that is $T$ close to f. Then there exists an absolute constant $C=C(T)>0$, independent of $\Delta$ from (12), so that
  • for each integer $k\geq k_{0}$, there is a unique local minimum $(r_{k},h_{1}(r_{k}))$ of $h_{1}$ that is $C$ close to the local minimum $(q_{k},f_{1}(q_{k}))$ of $f_{1}$, i.e. $r_{k}\in[q_{k}-C,q_{k}+C]$ and $h_{1}(r_{k})\in[f_{1}(q_{k})-C,f_{1}(q_{k})+C]$.
  • any other local minimum $(r,h_{1}(r))$ of $h_{1}$ satisfies $h_{1}(r)\geq-C$.
Proof.
Recall $f_{1}$ decays with slope $-1/n$ on intervals $(b_{k},q_{k})$, increases with slope $1/m$ on $(q_{k},c_{k})$ and is 00 outside these intervals. Since $\sup_{q>0}|f_{1}(q)-h_{1}(q)|\leq T$ we obviously have that $h_{1}(q)<-T$ implies $b_{k}<q<c_{k}$. Moreover, since $f_{2}(q)\geq 0$ everywhere and $|f_{2}(q)-h_{2}(q)|\leq T$ on $q\in(0,\infty)$, the graph of $h_{2}$ has no points with second coordinate less than $-T$. Since any local maximum of $h_{1}$ is a local minimum of $h_{2}$, it follows that there are no local maxima of $h_{1}$ with function value smaller than $-T$. Hence we have the following: On any interval $[b,c]$ where on the interior $(b,c)$ we have $h_{1}(q)<-T$ and at the boundary $h_{1}(b)=h_{1}(c)=-T$, the function $h_{1}(q)$ must be equal to some fixed trajectory $T_{\textbf{p},\textbf{q}}$. This means $h_{1}(q)$ first decreases starting from $b$ with slope $-1/n$ up to some local minimum attained at some $r>b$, and then starts increasing with slope $1/m$ until it reaches $c>r>b$ with $h_{1}(c)=-T$ again. We clearly have $[b,c]\subseteq[b_{k},c_{k}]$ for some $k$ as outside $f_{1}(q)=0$ and thus $h_{1}(q)\geq-T$, contradiction to $h_{1}(q)<-T$ in a right resp. left neighborhood of $b$ resp. $c$. More precisely, both $b,c$ must lie in one of the two subintervals (for now this may or may not be the same for $b$ and $c$)
(14) $[b_{k},b_{k}+2nT],\qquad[c_{k}-2mT,c_{k}]$
since in the remaining interval $(b_{k}+2nT,c_{k}-2mT)$ we have $f_{1}(q)<-2T$ and thus $h_{1}(q)<-T$, but $h_{1}(b)=h_{1}(c)=-T$.
We may assume the two subintervals in (14) are disjoint for all large $k$. Indeed, this is in particular true if
(15) $\lim_{k\to\infty}f_{1}(q_{k})=-\infty,$
as then by (13) we have $c_{k}-b_{k}=-(m+n)f_{1}(q_{k})\to\infty$. However (15) is in turn always true if $t_{k}^{n/m}\Phi(t_{k})\to 0$ as $k\to\infty$, as can be checked via (12). Now if otherwise $\Phi(t_{k})\gg t_{k}^{-n/m}$ on a subsequence of $k$, then by (1) it is easily seen that $Bad(\Phi)\supseteq Bad(t\to t^{-n/m})=Bad$ contains the usual set of badly approximable real $m\times n$ matrices (in fact our assumptions on $\Phi$ imply identity). Then the claim of Theorem 2.1 follows from the well-known result by Schmidt [11] (or alternatively derived from the variational principle) that the set of badly approximable matrices has full Hausdorff dimension $mn$.
From the disjointness, it is obvious from the above description of $h_{1}$ on $[b,c]$ that for each $k$ there can be at most one pair $b,c$ as above with the property
(16) $b_{k}\leq b\leq b_{k}+2nT,\qquad c_{k}-2mT\leq c\leq c_{k}.$
We consider $k$ fixed now and we write $r_{k}$ for the local minimum position $r$, which indeed will be the value stated in Lemma 4.3. In particular $|b-b_{k}|\ll_{T,m,n}1$ and $|c-c_{k}|\ll_{T,m,n}1$ uniformly in $k$.
For all other pairs where either both $b,c$ are in $[b_{k},b_{k}+2nT]$ or both are in $[c_{k}-2mT,c_{k}]$, since the slope of $h_{1}$ is bounded between $-1/n$ and $1/m$ and $h_{1}(b)=h_{1}(c)=-T$, it is clear that $h_{1}$ is uniformly bounded from below on $[b,c]$ in terms of $m,n,T$ only. In fact from intersecting the line containing $(b,-T)$ with slope $-1/n$ with the line containing $(c,-T)$ with slope $1/m$, recalling at $r=r_{k}$ is the local minimum of $h_{1}$ within $[b,c]$, we get the bound
$\min_{q\in[b,c]}h_{1}(q)=h_{1}(r)=-T-\frac{c-b}{m+n}\geq-T-\frac{2T\max\{m,n\}}{m+n}=:-C_{1}.$
Now we consider the potential pair as in (16). In fact this pair must exist for any large $k$ by (15). Indeed, for large $k$ we have $f_{1}(q_{k})<\min\{-2T,-C_{1}-T\}$ hence $h_{1}(q_{k})\leq f_{1}(q_{k})+T<\min\{-T,-C_{1}\}$ and we observed that only for pairs in (16) this may happen, for large $k$. Now for this pair in (16) we have
$\min_{q\in[b,c]}h_{1}(q)=h_{1}(r_{k})=-T-\frac{c-b}{m+n}=-T-\frac{c_{k}-b_{k}}{m+n}+\frac{E}{m+n}$
with $0\leq E\leq 2(m+n)T$, so by (13) we conclude
$\min_{q\in[b,c]}h_{1}(q)=h_{1}(r_{k})\in[-\frac{c_{k}-b_{k}}{m+n}-T,-\frac{c_{k}-b_{k}}{m+n}+T]=[f_{1}(q_{k})-T,f_{1}(q_{k})+T].$
In other words
(17) $|h_{1}(r_{k})-f_{1}(q_{k})|\leq T.$
Moreover, $r_{k}$ can be determined via solution of $-T-(r_{k}-b)/n=-T-(c-r_{k})/m$ as
$r_{k}=\frac{bm+cn}{m+n}.$
Similarly
$q_{k}=\frac{b_{k}m+c_{k}n}{m+n}.$
Hence by (16) we get
$|r_{k}-q_{k}|\leq\frac{4mnT}{m+n}=:C_{2}\ll_{m,n}T.$
So together with (17), indeed the points $(q_{k},f_{1}(q_{k}))$ and $(r_{k},h_{1}(r_{k}))$ are at most $C_{3}:=\max\{C_{2},T\}\ll_{m,n}T$ apart with respect to maximum norm on $\mathbb{R}^{2}$, proving the first claim of the lemma. The second is obvious as from the arguments above any other local minimum $\tilde{q}\notin\{r_{1},r_{2},\ldots\}$ of $h_{1}(q)$ satisfies $h_{1}(\tilde{q})\geq\min\{-T,-C_{1}\}=-C_{1}$, so we can put $C=\max\{C_{1},C_{3}\}$ to satisfy both claims. Note that $C$ is independent of $\Delta$ as requested. ∎
Given $\varepsilon>0$, start with $T=T_{0}(\varepsilon)$ induced by Theorem 3.1. With $C=C(T)$ from Lemma 4.3, take the square $[q_{k}-C,q_{k}+C]\times[f_{1}(q_{k})-C,f_{1}(q_{k})+C]$ and let $X_{k}\subseteq\mathbb{R}$ be the interval obtained from its projection to the second axis along the line with slope $-1/n$ and $Y_{k}\subseteq\mathbb{R}$ the interval induced by its projection to the second axis along the line with slope $1/m$. This means we identify $X_{k},Y_{k}$ with real intervals by dropping the first coordinate 00. Then by the formula of trajectories (9) and as at the minimum the two component functions within the maximum coincide, the integer vectors $(\textbf{p}_{k},\textbf{q}_{k})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\textbf{0}\})$ inducing $(r_{k},h_{1}(r_{k}))$, depending on h and thus the choice of $A$, satisfy
$\log\|\textbf{q}_{k}\|\in X_{k},\qquad\log\|A\textbf{q}_{k}-\textbf{p}_{k}\|\in Y_{k}.$
Since it is easy to see that $X_{k}\subseteq[f_{1}(q_{k})+q_{k}/n-D,f_{1}(q_{k})+q_{k}/n+D]$ and $Y_{k}\subseteq[f_{1}(q_{k})-q_{k}/m-D,f_{1}(q_{k})-q_{k}/m+D]$ for some absolute $D=D(C)>0$ independent of $k$ and $\Delta$, taking exponential map we see that uniformly in $k$
(18) $\|\textbf{q}_{k}\|\asymp_{D}e^{f_{1}(q_{k})+q_{k}/n}=t_{k},\qquad\|A\textbf{q}_{k}-\textbf{p}_{k}\|\asymp_{D}e^{f_{1}(q_{k})-q_{k}/m}=\frac{\Phi(t_{k})}{\Delta}.$
For the identities we have used (12), and the latter requires a short calculation. Recall $T$ is fixed and so are $C,D$, independent of $\Delta$. Hence, by property (C1), choosing $\Delta$ large enough we can guarantee that $\Phi(t_{k})/\Delta<K^{-1}\Phi(\|\textbf{q}_{k}\|)$ for all $k\geq 1$, where $K=K(D)=K(T)$ is the implied upper constant in the right formula of (18). Thus
(19) $\Phi(\|\textbf{q}_{k}\|)\ll_{D}\|A\textbf{q}_{k}-\textbf{p}_{k}\|<K\frac{\Phi(t_{k})}{\Delta}<K\cdot(K^{-1}\Phi(\|\textbf{q}_{k}\|))=\Phi(\|\textbf{q}_{k}\|).$
Again since $C$ and thus $D$ is fixed once $T$ is chosen, we may deduce that
$\Phi(\|\textbf{q}_{k}\|)\ll_{T}\|A\textbf{q}_{k}-\textbf{p}_{k}\|<\Phi(\|\textbf{q}_{k}\|).$
The right estimate shows that any $A$ whose combined graph is $T$ close to the template f lies in $W(\Phi)$. On the other hand, the left estimate shows that the $(\textbf{p}_{k},\textbf{q}_{k})\in\mathbb{Z}^{m}\times\mathbb{Z}^{n}$ inducing $(r_{k},h_{1}(r_{k}))$ do not induce considerably better approximations, so for $A$ to lie in all $W(\Phi/c)$, $c>1$ we may restrict to considering other integer vectors $(\textbf{p},\textbf{q})\in\mathbb{Z}^{m}\times\mathbb{Z}^{n}$. However, by Lemma 4.3 any $(\textbf{p},\textbf{q})\in\mathbb{Z}^{m}\times(\mathbb{Z}^{n}\setminus\{\textbf{0}\})$ not of the form $(\textbf{p}_{k},\textbf{q}_{k})$ induces a local minimum point $(r,h_{1}(r))$ with $h_{1}(r)\geq-C$, which by a short calculation and (1) means
$\|A\textbf{q}-\textbf{p}\|\gg_{C}\|\textbf{q}\|^{-n/m}\geq\Phi(\|\textbf{q}\|),$
and again since $C$ is fixed once $T$ is chosen also
$\|A\textbf{q}-\textbf{p}\|\gg_{T}\Phi(\|\textbf{q}\|).$
This argument shows that any $A$ with combined graph $T$ close to f cannot lie in all $W(\Phi/c)$, $c>1$. Combining our observations, indeed any such $A$ belongs to $Bad(\Phi)$.
4.3. Proof of Lemma 4.2
As announced, our proof is very similar to [6, Theorem 3.17]. We use the notation in [6, § 4.1]. Following [6, Definition 4.5], the local contraction rate of a template f at $q>0$ equals the cardinality of the set
(20) $\{(i_{+},i_{-}):i_{+}\in S_{+},i_{-}\in S_{-},\;\;i_{+}<i_{-}\},$
where $S_{-},S_{+}$ depend on $q$ and form a partition of $\{1,2,\ldots,m+n\}$, we omit stating the full intricate precise definition from [6, Definition 4.5]. In the next proposition we determine the contraction rates on $(0,\infty)$. Denote $Y_{\mathbb{Z}}=Y\cap\mathbb{Z}$ for a set $Y\subseteq\mathbb{R}$.
Proposition 4.4.
The local contraction rate of the template constructed in § 4.1 is given as
$\delta(q)=\begin{cases}mn,\qquad\qquad\text{if }\;q\notin\cup(b_{k},q_{k})\\ mn-m,\qquad\text{if }\;q\in\cup(b_{k},q_{k}).\end{cases}$
Proof.
In intervals of the form $(a_{k},b_{k})=(c_{k-1},b_{k})$, i.e. the interior of $I_{k}$, where all $f_{j}$ vanish, the only interval of equivalence as defined in [6, Definition 4.5] is $(0,mn]_{\mathbb{Z}}=\{1,2,\ldots,m+n\}$ and $S_{-}=\{m+1,m+2,\ldots,m+n\}$ and $S_{+}=\{1,2,\ldots,m\}$, so the cardinality of (20) is $mn$, as large as possible. On intervals $(b_{k},q_{k})$, the intervals of equivalence are $(0,1]_{\mathbb{Z}}=\{1\}$ and $(1,mn]_{\mathbb{Z}}=\{2,3,\ldots,m+n\}$ and $S_{-}=\{1,m+2,m+3,\ldots,m+n\}$ and $S_{+}=\{2,3,\ldots,m+1\}$, so the set (20) has cardinality $mn-m$. Finally, on intervals $(q_{k},c_{k})$, the intervals of equivalence are $(0,m+n-1]_{\mathbb{Z}}=\{1,2,\ldots,m+n-1\}$ and the singleton $(m+n-1,m+n]_{\mathbb{Z}}=\{m+n\}$, but then again $S_{-}=\{m+1,m+2,\ldots,m+n\}$ and $S_{+}=\{1,2,\ldots,m\}$. So the cardinality of (20) is again $mn$. ∎
By the proposition, it is clear that the local minima of the average contraction rate are taken at the values $q_{k}$, and the local maxima at values $b_{k}$, so that
[…]
From (10) and (12) if $\tau<\infty$ we see
(22) $q_{k}=\frac{\tau+1+o(1)}{m^{-1}+n^{-1}}\cdot\log t_{k},\qquad k\to\infty,$
and similarly we can just assume the first factor grows if $\tau=\infty$. In either case (11) implies
(23) $\lim_{k\to\infty}\frac{q_{k+1}}{q_{k}}=\infty.$
On the other hand, since by (9) and Dirichlet’s Theorem obviously $-1/n\leq f_{1}(q_{k})/q_{k}\leq 0$, relations (12), (13) imply $b_{k}\asymp c_{k}\asymp q_{k}$, with implied constants depending on $\Delta$. Combining with (23) yields $b_{k+1}/q_{k}\to\infty$ or $q_{k}/b_{k+1}\to 0$ as $k\to\infty$. Since the local contraction rate equals $mn$ on the long interval $(q_{k},b_{k+1})$, it is further easy to see that the right expression in (21) is $mn$, indeed
[…]
the reverse inequality being trivial. Moreover, by Proposition 4.4 and (23) the left expression in (21) can be estimated
[…]
$=\liminf_{k\to\infty}\frac{mn(b_{k}-q_{k-1})+(mn-m)(q_{k}-b_{k})}{q_{k}}$
$=mn\cdot(u_{k}-o(1))+(mn-m)(1-u_{k})$
$=mn(1-o(1))-(1-u_{k})m,$
as $k\to\infty$, where we have put
$u_{k}:=\frac{b_{k}}{q_{k}}.$
We determine the limit of $u_{k}$. The relations (12) and (22) imply
$f_{1}(q_{k})=\log t_{k}-q_{k}/n=q_{k}\cdot\left(\frac{m^{-1}+n^{-1}}{\tau+1}-\frac{1}{n}+o(1)\right),\quad k\to\infty,$
so further by (13), (22) as $k\to\infty$ we get
$b_{k}=q_{k}+nf_{1}(q_{k})=q_{k}\cdot\left(1+\frac{n/m+1}{\tau+1}-1+o(1)\right)=q_{k}\left(\frac{n/m+1}{\tau+1}+o(1)\right).$
Hence
$\lim_{k\to\infty}u_{k}=\frac{m+n}{m(\tau+1)}.$
Inserting in (24) and a short simplification indeed leads to the claimed lower bound. As observed below Theorem 2.1, the reverse upper bound is well-known.
4.4. Conclusion
Given $\varepsilon>0$, we have shown that for large $T$ the set of matrices $T$ close to the constructed template have Hausdorff and packing dimension at least $\underline{\delta}(\textbf{f})-\varepsilon$ and $\overline{\delta}(\textbf{f})-\varepsilon$, respectively. However, since the latter values are the predicted dimension of Theorem 2.1 and $\varepsilon$ can be arbitrarily small, the claimed lower bounds follow.
5. Proof of Theorem 2.1 for monotonic $\Phi$
5.1. The case $\tau<\infty$
Now drop the assumption (C1) and assume instead $\Phi$ is decreasing and $\tau<\infty$. We again use the construction of § 4.1 but will have to specify the involved lacunary sequence $t_{k}$. Analyzing the proof of Lemma 4.1, the only point where we needed (C1) was to conclude that for the constructed $\textbf{p}_{k},\textbf{q}_{k}$ with
$\|\textbf{q}_{k}\|\asymp t_{k},\qquad\|A\textbf{q}_{k}-\textbf{p}_{k}\|\asymp\Phi(t_{k})$
for any $\textbf{q}^{\prime}$ of comparable norm as $\textbf{q}_{k}$, we also have
$\Phi(\|\textbf{q}^{\prime}\|)\asymp\Phi(\|\textbf{q}\|).$
So to derive the same result for any non-increasing approximation function $\Phi$ with property (1), the key is again to prove existence of a sequence $t_{k}$ where this holds for similarly derived $\textbf{p}_{k},\textbf{q}_{k}$. This results in verifying the following rather elementary lemma on real functions, whose proof however is not as straightforward as it may appear at first sight.
Lemma 5.1.
Let $\Phi$ be decreasing and inducing $\tau<\infty$, and $c>1$ be fixed. Then there exists a sequence of integers $t_{k}$ satisfying (10) and (11) and so that additionally (C1) holds for any $t=t_{k}$, i.e. there is an explicitly computable fixed $d=d(c,\tau)>1$ independent of $k$ so that
(25) $d^{-1}\Phi(t_{k})\leq\Phi(\tilde{t})\leq d\Phi(t_{k})$
holds for any $k$ and any integer $\tilde{t}$ in the range
$c^{-1}t_{k}\leq\tilde{t}\leq ct_{k}.$
The proof of the sole implication (25) would be rather simple, however incorporating the additional condition (10) makes it slightly technical. In this context, we point out that (25) is in general not true for an arbitrary sequence $t_{k}$ satisfying (10) and (11) when we only assume $\Phi$ to be decreasing. Indeed, there may be problems when $\Phi$ decays fast close to $t_{k}$. Moreover (25) certainly fails irrespective of $d$ for all large $t_{k}$ if for example $\Phi(t)=e^{-t}$ inducing $\tau=\infty$, excluded in the lemma. Indeed we need a different argument in this case. We use both the monotonicity of $\Phi$ and $\tau<\infty$ to show that we are still able to find a suitable sequence $t_{k}$ avoiding these obstacles. As a side note we remark that for the sake of Lemma 5.1 only, we need not assume (1) on $\Phi$.
Proof.
To prove the lemma, complementing $\tau$, we define the upper order of $1/\Phi$ at infinity
$\omega:=\limsup_{t\to\infty}-\frac{\log\Phi(t)}{\log t}.$
For simplicity let us assume $\omega<\infty$, otherwise some minor twists in the proof below are required. Keep in mind that the monotonicity of $\Phi$ and $\omega<\infty$ easily imply for any $\sigma>1$ we have
(26) $\limsup_{N\to\infty}-\left(\frac{\log\Phi(\sigma N)}{\log(\sigma N)}-\frac{\log\Phi(N)}{\log N}\right)\leq 0,$
in particular
(27) $\limsup_{N\to\infty}-\left(\frac{\log\Phi(N+1)}{\log(N+1)}-\frac{\log\Phi(N)}{\log N}\right)\leq 0.$
Essentially this means that the function $t\mapsto-\log\Phi(t)/\log t$ can only decrease very slowly for large $t$ (whereas it may increase arbitrarily rapidly). We distinguish two cases.
Case 1: $\omega>\tau$, that is the upper order strictly exceeds the lower order of $1/\Phi$. Let $\delta_{k}\in(0,\omega-\tau)$ be a sequence with $\delta_{k}\to 0$ slowly. Fix $k$ for the moment. By choice of $\delta_{k}$ and definition of $\omega$, there exists an (arbitrarily large) integer $\ell=\ell_{k}$ satisfying $-\log\Phi(\ell_{k})/\log\ell_{k}>\tau+\delta_{k}$. Hence, by (27), if $\ell_{k}$ was chosen large enough, then by definition of $\tau$ there is some integer $r_{k}>\ell_{k}$ with $-\log\Phi(r_{k})/\log r_{k}$ between $\tau+\delta_{k}/2$ and $\tau+\delta_{k}$. Choose $r_{k}$ minimal with this property. Again by definition of $\tau$ at some point $s_{k}>r_{k}$ we have $-\log\Phi(s_{k})/\log s_{k}<\tau+\delta_{k}/3$. By (26) for some $\sigma\leq c^{2}$, we may assume $s_{k}$ is of the form $c^{2g}r_{k}$ for an integer $g=g(k)\geq 1$, otherwise we slightly enlarge it without affecting the estimates above. If we again choose $s_{k}$ minimal with this property and $\ell_{k}$ was chosen large enough, again by (26) we can assume conversely $-\log\Phi(s_{k})/\log s_{k}>\tau$. (Moreover it follows from montonicity of $\Phi$ that $s_{k}/r_{k}\to\infty$ as $k\to\infty$, however we will not explicitly need this). So by this argument we find a sequence of integer pairs $r_{k},s_{k}$ with
$r_{1}<s_{1}<r_{2}<s_{2}<\cdots$
and
(28) $\tau<-\frac{\log\Phi(s_{k})}{\log s_{k}}<\tau+\frac{\delta_{k}}{3}<\tau+\frac{\delta_{k}}{2}<-\frac{\log\Phi(r_{k})}{\log r_{k}}<\tau+\delta_{k}.$
We want to contradict this if the implication (25) fails for
$d=c^{2\omega}<\infty.$
For this, we split the interval $[r_{k},s_{k}]$ into $g=g(k)\geq 1$ adjacent subintervals of constant ratio $c^{2}$, i.e. of the form
$[r_{k},s_{k}]=[r_{k},c^{2}r_{k}]\cup[c^{2}r_{k},c^{4}r_{k}]\cup\cdots\cup[c^{2(g-1)}r_{k},c^{2g}r_{k}]$
Denote the subintervals by $I_{i}=[a_{i},b_{i}]$, for $1\leq i\leq g$, omitting $k$ in the notation, and formally let $a_{g+1}:=b_{g}=s_{k}$. Note that $b_{i}=a_{i+1}=c^{2}a_{i}$ for $1\leq i\leq g$ and
(29) $a_{i+1}=c^{2}a_{i},\qquad 1\leq i\leq g.$
Now assume contrary to (25) that on each subinterval $I_{i}$ we have
(30) $\frac{\max_{t\in I_{i}}\Phi(t)}{\min_{t\in I_{i}}\Phi(t)}=\frac{\Phi(a_{i})}{\Phi(a_{i+1})}>d=c^{2\omega},\qquad 1\leq i\leq g.$
We may assume $r_{1}=a_{1}(1)\leq a_{1}(k)=r_{k}$ was chosen large enough that
(31) $\Phi(t)>t^{-2\omega},\qquad t\geq r_{1}.$
Combination of the above estimates (29), (30), (31) readily implies
$-\frac{\log\Phi(r_{k})}{\log r_{k}}=-\frac{\log\Phi(a_{1})}{\log a_{1}}<-\frac{\log\Phi(a_{2})}{\log a_{2}}<\cdots<-\frac{\log\Phi(a_{g+1})}{\log a_{g+1}}=-\frac{\log\Phi(s_{k})}{\log s_{k}}$
contradiction to (28). So for each $k$ there must be some interval $I_{j}$, $j=j(k)$, where the claim (25) holds. Moreover, by (28) we have
$-\frac{\log\Phi(t)}{\log t}\in[\tau,\tau+\delta_{k}],\qquad t\in I_{j}$
so since $\delta_{k}\to 0$ this tends to $\tau$. So if we let $t_{k}=ca_{j}$, $j=j(k)$ as above such that $[c^{-1}t_{k},ct_{k}]=I_{j}$, and increase $\ell_{k}$ and thus $t_{k}$ fast enough in each step, then indeed all claimed properties are satisfied. Finally the case $\omega=\infty$ requires just minor modifications in the argument, we leave the details to the reader.
Case 2: $\omega=\tau$. Then clearly it suffices to find any lacunary sequence $t_{k}$ with property (25). Start with large integer $\ell_{1}>0$. If $\Phi(c^{2}\ell_{1})>c^{-4\tau}\Phi(\ell_{1})$ then by monotonicity $c^{-4\tau}\Phi(\ell_{1})\leq\Phi(t)\leq\Phi(\ell_{1})$ for each $t\in[\ell_{1},c^{2}\ell_{1}]$, in other words
$c^{-4\tau}\leq\frac{\Phi(t)}{\Phi(\ell_{1})}\leq\frac{\Phi(t)}{\Phi(c\ell_{1})}\leq 1$
for such $t$. Note that $t$ lies in $[c^{-1}(c\ell_{1}),c(c\ell_{1})]$, hence with $c$ and $t_{1}=c\ell_{1}$ as in the theorem.
Hence we may put $t_{1}=c\ell_{1}$ and are done with first step for implied constant $d=c^{4\tau}$, and then we choose $\ell_{2}>\ell_{1}$ large, inducing $t_{2}>t_{1}$ with the same argument. So assume opposite $\Phi(c^{2}\ell_{1})\leq c^{-4\tau}\Phi(\ell_{1})$. Then take $\widetilde{\ell_{1}}=c^{2}\ell_{1}$ instead. Again by the same argument we can take $t_{1}=c\widetilde{\ell_{1}}=c(c^{2}\ell_{1})=c^{3}\ell_{1}$ unless if $\Phi(c^{2}\widetilde{\ell_{1}})\leq c^{-4\tau}\Phi(\widetilde{\ell_{1}})$. We repeat this process and see that the argument to find $t_{1}$ for the constant $d=c^{4\tau}$ can only fail if for any integer $k\geq 1$ we have $\Phi(c^{2k}\ell_{1})\leq c^{-4\tau k}\Phi(\ell_{1})$. However as $k\to\infty$ this would imply the upper order of $\Phi$ is at least $\omega\geq 4\tau k/(2k)=2\tau>\tau$, contradiction to our assumption of Case 2. ∎
With the lemma at hand, it is easy to conclude. By Theorem 3.1, for any $\varepsilon>0$ and some $T=T_{0}(\varepsilon)$ large enough we have uniformly in all templates $\boldsymbol{f}$
(32) $\dim_{H}(\{A:\sup_{q>0}\max_{1\leq j\leq m+n}|f_{j}(q)-h_{j}(q)|\leq T\})\geq\underline{\delta}(\boldsymbol{f})-\varepsilon,$
where $h_{j}=h_{j,A}$ are the component functions of the combined graph associated to $A$. Now starting with decreasing $\Phi$ satisfying (1) and $\tau<\infty$, we may apply Lemma 5.1 for any $c>0$ which equips us with some induced lacunary sequence $t_{k}=t_{k}(c)$ from which we construct the template $\textbf{f}=\textbf{f}_{c}$ as in § 4.1. Now we can copy precisely the proof of Lemma 4.1 for $T$ above, with the sole twist that we use Lemma 5.1 in place of property (C1) to obtain (19). Since any $A$ within the left hand side set of (32) belongs to $Bad(\Phi)$ again and $\varepsilon>0$ is arbitrarily small, the Hausdorff dimension claim follows with Lemma 4.2 again which still holds for the same reasons. Similar arguments apply for packing dimension.
Remark 2.
It is possible to obtain the final deduction using a non-uniform version Theorem 3.1. However, this requires to first prove a uniform variant of Lemma 5.1, which is more technical.
5.2. The case $\tau=\infty$
Since as observed above in this case our key Lemma 5.1 above fails, we use a different argument. Let us denote for clarity the set $Exact(\Phi)=Exact^{m,n}(\Phi)$ for given $m,n$, and similarly define $W^{m,n}(\Phi)$. First recall that it then follows from [10] that for $n=1$ the set $Exact^{m,1}(\Phi)\subseteq\mathbb{R}^{m}$ has full packing dimension $m$, in particular it is non-empty (since $\Phi$ is monotonic here, the latter weaker claim also follows directly from Jarník [7, Satz 6]).
Using essentially the method by Moshchevitin [9, Theorem 12] we show the following.
Lemma 5.2.
Let $\Phi$ induce $\tau(\Phi)>\theta$. Let $\textbf{a}\in Exact^{m,1}(\Phi)\subseteq\mathbb{R}^{m}$ be a column vector. Then for almost all matrices $U\in\mathbb{R}^{m\times(n-1)}$ with respect to $(n-1)m$-dimensional Lebesgue measure, the composed matrix $A:=(\textbf{a},U)\in\mathbb{R}^{m\times n}$ with first column a followed by $U$ on the right also lies in $Exact^{m,n}(\Phi)\subseteq\mathbb{R}^{m\times n}$.
Remark 3.
The claim of the lemma applies, upon modifying $\theta$, more generally to extensions of arbitrary matrices (in place of a column vector) via addition of columns, by the same argument. Moreover we may replace $Exact^{.,.}(\Phi)$ by $Bad^{.,.}(\Phi)$ throughout.
Proof.
Note that for any matrix $U$ as in the lemma, for any $q\in\mathbb{Z},\textbf{p}\in\mathbb{Z}^{m}$ if we let $\textbf{q}=(q,0,0,\ldots,0)\in\mathbb{Z}^{n}$ we have
(33) $\|\textbf{q}\|=|q|,\qquad\|A\textbf{q}-\textbf{p}\|=\|\textbf{a}q-\textbf{p}\|.$
In particular $\textbf{a}\in Exact^{m,1}(\Phi)\subseteq W^{m,1}(\Phi)$ implies $A\in W^{m,n}(\Phi)$. The substantial part is to show conversely that $A\notin W^{m,n}(\Phi)/c$ for any $c>1$ and Lebesgue almost all $U$. Note that $\textbf{a}\notin W^{m,1}(\Phi)/c$ by assumption. For this we use the method in [9]. Let $c>1$ fixed. For $\textbf{p}\in\mathbb{Z}^{m},\textbf{q}\in\mathbb{Z}^{n}$ and writing $\textbf{a}=(a_{1},\ldots,a_{m})$, define for $1\leq j\leq m$ the sets
$J_{N}(p_{j},q_{1})=(-p_{j}+q_{1}a_{j}-\Phi(N)/c,-p_{j}+q_{1}a_{j})\subseteq\mathbb{R},\qquad N=1,2,\ldots.$
Derive
$\Omega_{N}(\textbf{p},\textbf{q})=\{U\in\mathbb{R}^{m(n-1)}:q_{2}u_{j,1}+\cdots+q_{n}u_{j,n-1}\in J_{N}(p_{j},q_{1}):1\leq j\leq m\}$
where $U=(u_{i,k})$ for $11\leq i\leq m,1\leq k\leq n-$ 1, are the entries of $U$. Let further
$\Omega_{N}=\bigcup_{\textbf{p},\textbf{q}:\;\|\textbf{q}\|=N,\;\textbf{q}^{\ast}\neq 0}\Omega_{N}(\textbf{p},\textbf{q})$
where $\textbf{q}^{\ast}=(q_{2},\ldots,q_{n})$, so the union is taken over integer vectors where not all $q_{2},\ldots,q_{n}$ vanish and with maximum of $|q_{1}|,\ldots,|q_{n}|$ equal to $N$. Note that we do not need to take into account vectors with $\textbf{q}^{\ast}=0$ since $\textbf{a}\notin W^{m,1}(\Phi)/c$ and (33). Then if converse to our assumption $A\in W^{m,n}(\Phi)/c$ for $A=(\textbf{a},U)$, then the matrix $U$ lies in the limsup set of the $\Omega_{N}$, that is it belongs to infinitely many $\Omega_{N}$. But for this to be true for a positive measure set of matrices $U$, by Borel-Cantelli lemma we require
$\sum_{N=1}^{\infty}\mu(\Omega_{N})=\infty,$
[…]
$\{(x_{1},\ldots,x_{n-1}):q_{2}x_{1}+\cdots+q_{n}x_{n-1}=-p_{j}+q_{1}a_{1}\}\subseteq\mathbb{R}^{n-1},\qquad 1\leq j\leq m,$
is at most $(\Phi(N)/c)\cdot(q_{2}^{2}+\cdots+q_{n}^{2})^{-1/2}\ll\Phi(N)\|\textbf{q}^{\ast}\|^{-1}$, so the body in $\mathbb{R}^{m(n-1)}$ where this is true for $1\leq j\leq m$ has volume $\ll\Phi(N)^{m}\|\textbf{q}^{\ast}\|^{-m}$. Moreover we can restrict to $\|\textbf{p}\|\ll N$ as otherwise $\Omega_{N}(\textbf{p},\textbf{q})=\emptyset$. Thus similar to [9] we calculate
$\mu(\Omega_{N})$ $\ll\Phi(N)^{m}N^{m}\sum_{q_{1}\in\mathbb{Z}:|q_{1}|\leq N}\sum_{q_{2},\ldots,q_{n}\in\mathbb{Z}:1\leq\max|q_{i}|\leq N}\frac{1}{\max_{2\leq i\leq n}|q_{i}|^{m}}$
$\ll\Phi(N)^{m}N^{m+1}\sum_{1\leq r\leq N}r^{n-m-2}$
$\ll\Phi(N)^{m}N^{\max\{1+m,n\}}(\log N)^{\delta(n,m+1)}$
where $\delta(a,b)=1$ if $a=b$ and $\delta(a,b)=0$ otherwise. Now since $\tau>\theta$ we know $\Phi(N)<N^{-\theta-\epsilon}$ for some $\epsilon>0$ and all large $N\geq N_{0}$. Hence by definition of $\theta$ the tail of the sum is at most $\sum_{N\geq N_{0}}N^{-1-m\epsilon}\log N$ which indeed converges. ∎
We have noticed above that $\textbf{a}\in Exact^{m,1}(\Phi)\subseteq\mathbb{R}^{m}$ exist. Taking any such a, for any $A$ derived as in the lemma $A\in Exact^{m,n}(\Phi)\subseteq Bad^{m,n}(\Phi)$ as well. The Hausdorff dimension of the set of such matrices $A$ is by construction at least $(n-1)m$, for as $W_{\infty}$, so the Hausdorff dimension claim (4) is proved. For the packing dimension result (5) we combine Lemma 5.2 with the estimate
(34) $\dim_{P}(X\times Y)\geq\dim_{P}(X)+\dim_{H}(Y)$
for any measurable $X,Y$ due to Tricot [13] with $X=Exact^{m,1}(\Phi)\subseteq\mathbb{R}^{m}$ and $Y$ the set of $U$ from Lemma 5.2. Then the claim follows from aforementioned full packing dimension of $X$. The proof of the case $\tau=\infty$ is complete as well.
6. Proof of Theorem 2.3
6.1. The proof
If $\tau=\infty$ both (a), (b) follow from the proof in § 5.2, so assume $\tau<\infty$. Again we proceed similar to § 5.2. However, for (a), instead of using [10], we start directly with the set of real vectors a in $\mathbb{R}^{m\times 1}$ in $Exact^{m,1}(\Phi)$ of Hausdorff dimension $(m+1)/(\tau+1)>0$ derived in [1], which applies since $\tau>\theta\geq(n+1)/m>1/m$ and $\Phi$ is assumed decreasing. Then since $\tau>\theta$ again Lemma 5.2 applies as before and we can add $(n-1)m$ by the property
(35) $\dim_{H}(X\times Y)\geq\dim_{H}(X)+\dim_{H}(Y)$
for any measurable $X,Y$, see Tricot [13] again. For claim (b), for any $\Phi$ inducing large enough $\tau$ (not necessarily monotonic), where $\tau>(5+\sqrt{17})/2$ suffices, the full packing dimension result from [10] used in § 5.2 in the vector setting $n=1$ still applies. Hence if $\tau>\theta$ as well, again by Lemma 5.2 we get (7), (8), where for the latter we also use (34).
6.2. Verification of Remark 1
We finally verify the claim in Remark 1. By [5], since $\tau>\theta>1$ and ()($\ast$)( ∗ ) holds for $\Phi$, again for $m=n=1$ the set $Exact^{1,1}(\Phi)\subseteq\mathbb{R}$ has full Hausdorff dimension $2/(1+\tau)$. Now take any real number $a\in Exact^{1,1}(\Phi)$ and define the first column of $A$ as $\textbf{a}=(a,a,\ldots,a)\in\mathbb{R}^{m\times 1}$. It is easy to see that similarly $\textbf{a}\in Exact^{m,1}(\Phi)$. Indeed, if $(q,p)$ is an integer vector realizing a good approximation $p/q$ to $a$, then $(q,p,p,\ldots,p)$ yields the same approximation quality $|qa-p|=\|q\textbf{a}-(p,\ldots,p)\|$ by choice of maximum norms, hence $\textbf{a}\in W^{m,1}(\Phi)$ is implied by $a\in W^{1,1}(\Phi)$. Conversely if for $(q,p_{1},\ldots,p_{m})$ we have $\|q\textbf{a}-\textbf{p}\|<\Phi(q)/c$ for some $c>1$, then also $|qa-p_{1}|<\Phi(q)/c$, which shows that if $\textbf{a}\in W^{m,1}(\Phi)/c$ then also $a\in W^{1,1}(\Phi)/c$. Since this is not true for any $c>1$ as $a\in Exact^{1,1}(\Phi)$, it cannot be true for any $c>1$ for a either. Finally by $\tau>\theta$ using our a above we can conclude with Lemma 5.2 and (35) again. We should notice however that the components of our constructed real matrices are not $\mathbb{Q}$-linearly independent.